# Camera-based 3D Semantic Scene Completion with Sparse Guidance Network

### v1

@@ Appendix A Appendix
Table 8: Number of model dimensions and depth of MSSD. Memory denotes training memory
Frames IoU (%) mIoU (%) Memory (G)
1 43.60 14.55 14.21
2 44.39 15.36 14.71
3 45.50 15.75 15.19
4 45.84 15.56 15.67
5 46.21 15.32 15.83
6 45.65 15.29 16.59
7 45.55 15.34 17.97
Table 9: Effect of temporal frames.
Figure 5: Visual comparison of our SGN-T with state-of-the-art methods on SemanticKITTI validation. Compared to VoxFormer-T and MonoScene, our SGN-T generates more precise segmentation boundaries . (labeled in red circles).
Efficiency analysis.
@@
We perform runtime experiments on a single V100 GPU. The mean value over the SemanticKITTI test set is reported. Our SGN-S, SGN-L, and SGN-T run in 327.71 ms, 315.35 ms, and 436.24 ms, respectively. We also tested the recent VoxFormer-T (261.46 ms) and OccFormer (322.87 ms) on the same platform with the officially provided weights for a fair comparison. Compared with these methods, our light version SGN-L has comparable latency but better applicability due to its lightweight (12.5M parameters) and less training memory consumption (7.16 G).
Qualitative Visualizations
We provide the visualization results of the proposed SGN-T on SemanticKITTI validation in Fig. 5. Compared to VoxFormer-T and MonoScene, our SGN-T generates more precise segmentation boundaries, especially on “plane” classes and large objects such as cars. Besides, SGN-T predicts more accurate SSC results and preserves more regional details in the short-range areas than other methods. For example, there are some wrong semantics and missing objects for VoxFormer-T and MonoScene in the short-range areas, as shown in the first and third rows in Fig. 5. However, we noticed that our SGN-T also missed some distant objects that are very small in RGB images. We explain that our SGN uses image features of the 1/16 scale, which may degrade the performance of objects in distant areas.

### v2

@@ IV-E Ablation studies
Impact of different training modes
To explore the effect of different training modes, i.e., two-stage and one-stage training, we provide the detailed comparison results with VoxFormer in Table VI. Line 3 of Table VI presents the results of our SGN-S with the two-stage training strategy. Note that when equipping SGN-S with a two-stage approach, the first stage remains the same as VoxFormer, and the parameters and training memory of the second stage are calculated for a fair comparison. In the same training configuration, Our SGN-S surpasses VoxFormer-S by a large margin regarding mIoU scores (+2.58% points). Even our one-stage SGN-S outperforms two-stage VoxFormer-S by 2.2% points in mIoU. It is worth noting that the model parameters of our SGN-S are only about half of those of VoxFormer-S. For the temporal version, our SGN-T achieves 46.21 IoU and 15.32 mIoU scores, boosting VoxFormer-T by 2.06% points in IoU and 1.97% points in mIoU. Our lightweight version, SGN-L, with only 12.5M parameters, also outperforms VoxFormer-T on mIoU and IoU scores while requiring only about half the training memory (7.16 G).
Abaltion on depth estimator
Our sparse voxel proposal network produces the seed voxels based on the depth map predicted by the depth estimator. The generated depth map contains 3D structure information, such as volume surfaces, which has a direct impact on the seed voxel proposal. To quantitatively analyze the impact of the depth estimator, we compare our SGN equipped with the monocular-based AdaBins [60] and stereo-based MobileStereoNet [53] with VoxFormer [12]. The results are presented in Table VII and show that using the stereo-based depth estimation brought significant performance improvements for both VoxFormer and our SGN, which means a stronger depth estimator that produces more accurate depth maps can boost the performance further. Notably, in the same configurations, our SGN consistently surpasses VoxFormer in all different ranges, demonstrating the effectiveness and superiority of our approaches.
Impact of image features
The 2D features provide a foundation for informative voxel features. We do detailed experiments in Table IX to explore the impact of the feature scale and image backbone. We see that using 2D features at a scale of 1/16 in ResNet 50 achieves the best IoU score and has comparable mIoU to other variants. And it strikes a balance between performance and model size. The results of using ResNet18 as the backbone are presented in the last line of Table IX and show that a more lightweight image backbone reduces the model parameters by 12.43M while the performance of the mIoU score drops by 0.47% points.
Figure 5: Effect of temporal frames. The frames are sampled every three frames. Memory denotes training memory.
Ablation on view transformation
The view transformation generates the initial 3D features for the subsequent hybrid guidance and informative voxel features. We further investigate the effect of different view transformation modules. We implement three commonly used modules, i.e., FLoSP in Monoscene [7], LSS [11], and cross-attention adopted from VoxFormer [12]. The results are presented in Table VIII, showing that using FLoSP contains fewer parameters while achieving comparable performance to other variants on both mIoU and IoU scores. Therefore, our SGN lifts the 2D features to 3D volume with the view transformation designed in the spirit of FLoSP.
Exploration on the threshold for seed voxels
We change the value $\theta$ to investigate the impact of different thresholds for selecting seed voxels. We calculate the average occupancy rate of seed voxels, mIoU score, and IoU score for variants with $\theta$ from 0.1 to 0.9. The results are shown in Figure 7, showing that the performance of the model first increases and then decreases as $\theta$ increases. And when $\theta=0.4$, the model achieves the best performance on both mIoU and IoU scores. Interestingly, we found our model still achieves notable mIoU and IoU scores when the ratio of the seed voxels is very low ($<5\%$ points). It shows that seed voxels with high confidence play an essential role in our semantic propagation.
Figure 6: Qualitative results of our SGN-S, SGN-L, and SGN-T on SemanticKITTI hidden test set. Our method can provide accurate semantic occupancy prediction and handle typical driving scenes such as crowded cars, shadows, tiny poles, and crossroads.
Figure 7: Impact of threshold values for seed voxels. The performance of the model first increases and then decreases as $\theta$ increases. And when $\theta=0.4$, the model achieves the best performance on both mIoU and IoU scores.
Number of model dimensions
The impact of the number of model dimensions of our SGN-S is evaluated and presented in Table X. The results reveal that using a large number of feature channels for 3D features boosts the models’ performance while increasing the model complexity. For example, the model with 64 dimensions contains fewer parameters and requires less training memory, although its performance drops by 0.28% points in IoU and 0.35% points in mIoU (Line3 vs Line4). We also provide the results of different model depths for the multi-scale semantic propagation module. We see that using different depths has comparable mIoU and IoU, demonstrating that our model with hybrid guidance and semantic propagation avoids the dependency on the heavy 3D model for processing 3D features.
Effect of temporal input.
Finally, we explore the impact of the number of temporal frames in Figure 5. We take historical frames to form the temporal input. As Figure 5 shows, the model’s performance on IoU scores first increases with the number of frames and then decreases. We explain that the camera extrinsic matrix from the historical frame to the current system is not always accurate, especially when the temporal interval is long. Therefore, the 3D points may project on the wrong image patches of the historical frames, which disturbs the learning of voxel features. To better balance the performance and memory consumption, our temporal version takes five frames (four past frames and the current frame).
IV-F Efficiency analysis.
@@
We perform runtime experiments on a single V100 GPU. The mean value over the SemanticKITTI test set is reported. Our SGN-S, SGN-L, and SGN-T run in 327.71 ms, 315.35 ms, and 436.24 ms, respectively. We also tested the recent VoxFormer-T (261.46 ms, $\sim$60M parameters) and OccFormer (322.87 ms, $\sim$200M parameters) on the same platform with the officially provided weights for a fair comparison. Compared with these methods, our lightweight version SGN-L achieves better mIoU and IoU scores and comparable latency, but with better applicability due to its lightweight (12.5M parameters) and less training memory consumption (7.16 G).
[…]
Figure 8: Qualitative results of our SGN-S on NYUv2 test set. Our method can provide accurate semantic occupancy prediction, demonstrating the generalization ability on the indoor scenes.
TABLE XI: Semantic Scene completion on NYUv2 test set. These compared methods are copy from Monoscene [7] and NDC-scene [32]. Bold and underline denote the best performance and the second-best performance, respectively.
Method IoU
$\blacksquare$ ceiling (1.37%)
$\blacksquare$ floor (17.58%)
$\blacksquare$ wall (15.26%)
$\blacksquare$ window (1.99%)
$\blacksquare$ chair (3.01%)
$\blacksquare$ bed (7.08%)
$\blacksquare$ sofa (4.70%)
$\blacksquare$ table (4.31%)
$\blacksquare$ tvs (0.47%)
$\blacksquare$ furniture (30.04%)
$\blacksquare$ objects (14.19%)
mIoU
[…]
MonoScene [7] 42.51 8.89 93.50 12.06 12.57 13.72 48.19 36.11 15.13 15.22 27.96 12.94 26.94
NDC-Scene [32] 44.17 12.02 93.51 13.11 13.77 15.83 49.57 39.87 17.17 24.57 31.00 14.96 29.03
SGN-S (ours) 44.85 14.67 93.56 13.36 10.35 14.64 48.59 37.47 14.08 15.75 30.31 12.07 27.71
IV-G Generalization on indoor scenes.
Although our proposed SGN mainly focuses on the outdoor driving scene as mentioned in Section I, we further provide detailed experiments on the NYUv2 [58] dataset to demonstrate the generalization ability on indoor scenes. NYUv2 comprises 1449 indoor scenes, represented as $240\times 144\times 240$ voxel grids labeled with 13 classes (11 semantics, 1 free, 1 unknown). The input resolution is $640\times 480$. Following [7, 32], we utilize a train/test splits of 795/654 scenes to perform evaluations on the test set at the scale of 1:4. Consistent with MonoScene [7], to verify the effectiveness on the indoor scenes, we utilize our single-image version SGN-S with the pre-trained EfficientNetB7 [62] as the image encoder and change the size $X\times Y\times Z$ to $60\times 36\times 60$. The 2D feature maps with 1/8 of the input resolution are taken for the subsequent processing. We apply [63] to generate monocular depth prediction and train SGN-S for 30 epochs using the AdamW [59] optimizer with the initial learning rate of 2e-4 and a weight decay of 1e-3.
As shown in Table XI, without any bells and whistles, our SGN-S achieves the best performance in terms of IoU score and delivers comparable results in mIoU score, which demonstrates the generalization ability of our proposed method on indoor scenes. For instance, SGN-S surpasses MonoScene by 0.77% and 2.34% points in mIoU and IoU scores, respectively. However, we observe that NDC-Scene outperforms SGN-S in mIoU score. We attribute this to the complexity and sensitivity of indoor scenes to the accuracy of the depth estimator. Additionally, SGN-S uses an image scale of only 1/8, which may impact the segmentation details, resulting in worse performance on some categories such as “TVs” and “objects,” as shown in Table XI. We believe that incorporating a more accurate depth estimator, more effective multi-scale feature fusion, and advanced loss designs, such as the frustum proportion loss used in [7], could further enhance performance. This will be a focus of our future work. We also present the qualitative results of our method and recent methods on the NYUv2 test set. As illustrated in Figure 8, even without a special design for the indoor scenarios, our SGN-S still generates precise semantic scene completion prediction, which further demonstrates the generalization ability of our method.
V Conclusion
This work focuses on camera-based semantic scene completion (SSC). Existing methods usually rely on sophisticated 3D models to directly process the coarse lifted 3D features that are not discriminative enough for clear segmentation boundaries. Therefore, we propose the one-stage SGN to propagate semantics from the semantic-aware seed voxels to the whole scene based on spatial geometry information. We first redesign the sparse voxel proposal network with the coarse-to-fine paradigm for dynamically and accurately selecting seed voxels. Then, we design hybrid guidance and effective voxel aggregation to enhance the intra-category feature separations and expedite the convergence of semantic propagation. Finally, the multi-scale semantic propagation is proposed for the final semantic scene completion. Extensive experiments on the SemanticKITTI and SSCBench-KITTI-360 benchmarks demonstrate the effectiveness of Our SGN, which achieves state-of-the-art performance while being more lightweight.
We hope our work can promote the exploration of model optimization and lightweighting in 3D scene understanding and provide innovative solutions for applications in scenarios with limited resources.